# D-tensor paraproducts and its caricatures

### v1

@@ Abstract
We generalize the $2$-tensor paraproduct decomposition result of [1] to $d$-tensors and expound on its caricatures. In particular, we show that for $A\in C^{d},f\in\Lambda^{\alpha}([0,1]^{d})$, $A(f)$ has an approximation $\tilde{A}(f)=(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f))$ for a fixed sequence of scales, $\mathbf{j}=(j_{1},j_{2},\ldots,j_{d})$, and the series $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f)$ is a taylor expansion of $A$. Additionally, we show the sequence of operators $(\tilde{\mathbf{v}}^{1}(f),\tilde{\mathbf{v}}^{2}(f),\ldots,\tilde{\mathbf{v}}^{d}(f))$ form a basis for the subspace comprised of linear functionals of the form $A(f)$, and the residual, $\Delta(A,f)=\tilde{A}(f)-A(f)\in\Lambda^{2\alpha}([0,1]^{d})$. Consequentially, we show one can obtain a Calderon-Zygmund type decomposition, $\Delta(A,f)_{\lambda}+\tilde{A}(f)_{\lambda}$ with the obtained paraproduct decomposition $A(f)=\Delta(A,f)+\tilde{A}(f)$. Our theoretical findings are supported by computational examples for d=2,3.
1 Introduction
We generalize the 2-tensor paraproduct decomposition established in [1] to $d$-tensors. A version of J.M. Bony’s celebrated result [2] was formulated in [1] for smooth compositions of mixed-holder functions which appear in a variety of matrix processing tasks [3, 4]. We extend this result for Holder functions of the form $A(f),f:[0,1]^{d}\to\mathbb{R},A\in C^{d}$ to obtain the paraproduct decomposition, $A(f)=\tilde{A}(f)+\Delta(A,f)$. Examination of the organs of $\tilde{A}(f)+\Delta(A,f)$ leads to several caricatures which are highlighted herein. Namely, in higher dimensions we show such a decomposition is a taylor expansion of $A$ on the image of $f$, the operators in the taylor expansion of $f$ form a basis for the subspace comprised of functions of the form $A(f)$, and the paraproduct decompostion, $A(f)=\tilde{A}(f)+\Delta(A,f)$, can readily be modified to obtain the Calderon-Zygmund decomposition, $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$.
Our theoretical findings are supported by computational examples of commonly encountered mathematical objects for $d=2,3$. For $d=2$ we consider the potential kernel, $A_{1}(f)=\frac{1}{f(x,y)}$ and the heat kernel $A_{2}(f)=e^{-\frac{f(x,y)}{\epsilon}}$ where $f(x,y)=\lVert x-y\rVert_{l_{2}}$. (we first organize $f(x,y)$ to be $\Lambda^{\alpha}([0,1]^{2})$ using the algorithm from [5]). Next we consider a cone singularity in the complex plane modulated by a graphic equalizer: $A_{3}(f)=\text{sin}(\frac{j_{x}+j_{iy}}{2}f(z))$. Subsequent processing of the low and high frequency components separately with Haar for $A_{1}(f),A_{2}(f)$, for example, leads to tensor compression suggesting similar algorithms to [6] could be explored for $d=2,3$ with such a decomposition. Furthermore, we show the time-varying singularity in the complex plane ($A_{3}(f)$) can be detected in the presence of the graphic equalizer. First, we restate the decomposition theorem for $2$-tensors:
Theorem 1.
Suppose $A\in C^{2}$, $f\in\Lambda^{\alpha}([0,1]^{2}),0<\alpha<\frac{1}{2}$, then we obtain the following discrete multiscale tensor decomposition of $A(f)$:
$A(f)=\sum_{j^{\prime}=0^{\prime}}^{N^{\prime}}\sum_{j=0}^{N}A^{\prime}(P^{j}P^{\prime j^{\prime}}(f))[P^{j+1}P^{\prime j^{\prime}+1}(f)-P^{j}P^{\prime j^{\prime}+1}(f)-P^{j+1}P^{\prime j^{\prime}}(f)+P^{j}P^{\prime j^{\prime}}(f)]$
$+A^{\prime\prime}(P^{j}P^{\prime j^{\prime}}(f))[P^{j+1}P^{\prime j^{\prime}}(f)-P^{j}P^{\prime j^{\prime}}(f)][P^{j}P^{\prime j^{\prime}+1}(f)-P^{j}P^{\prime j^{\prime}}(f)]+\Delta_{N,N^{\prime}}(A,f)$ (1)
$A(f)=\tilde{A}(f)+\Delta_{N,N^{\prime}}(A,f)$ (2)
@@ 1 Introduction
$\Delta_{N,N^{\prime}}(A,f)=A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{2})$ (3)
where $\Delta_{N,N^{\prime}}(A,f)$ is the residual between $A(f)$ and $\tilde{A}(f)$ up to scales $N,N^{\prime}$. $P^{j}P^{\prime j^{\prime}}$ is a convolution operators acting on $f$ with the tensor product of the scaling function, $\phi^{j,j^{\prime}}_{k,k^{\prime}}(x,x^{\prime})$ (here defined to be the characteristic function) at scales $j,j^{\prime}$ as its kernel. $R^{j,j^{\prime}}_{k,k^{\prime}}\coloneq I^{j}_{k}\times I^{j^{\prime}}_{k^{\prime}}$ is a dyadic rectangle constructed from taking the product of dyadic intervals $I^{j}_{k}\coloneq(2^{-j}k,2^{-j}(k+1)],k,j\in\mathbb{N}$ and $I^{j^{\prime}}_{k^{\prime}}\coloneq(2^{-j^{\prime}}k,2^{-j^{\prime}}(k^{\prime}+1)],k^{\prime},j^{\prime}\in\mathbb{N}$ such that $|R^{j,j^{\prime}}_{k,k^{\prime}}|=2^{-(j+j^{\prime})}$
This decomposition utilizes the connection between the finite difference of the scales of the kernels (here the kernels are the scaling functions) of the averaging operators acting on mixed holder functions to the fundamental theorem of calculus. We can use the same connection to establish the result in $d$-dimensions. We begin by defining mixed-Holder regularity for a $d$-dimensional function as well as the $d$-dimensional scaling function:
Definition 1.
The 1D scaling function is defined as:
$\phi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\phi(2^{-j}x-k),j,k\in\mathbb{N}$ (4)
where the scaling function, $\phi(x)$, is the characteristic function:
$\phi(x)\coloneq\begin{cases}1&\text{if }x\in(0,1]\\ 0&\text{if }x\notin(0,1]\end{cases}$ (5)
Definition 2.
The 1D wavelet function is defined as:
$\psi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\psi(2^{-j}x-k),j,k\in\mathbb{N}$ (6)
where the wavelet function, $\phi(x)$ is the Haar function:
$\psi(x)\coloneq\begin{cases}1&\text{if }0<x\leq\frac{1}{2}\\ -1&\text{if }\frac{1}{2}\leq x\leq 1\end{cases}$ (7)
Definition 3.
@@
$\phi(x_{1},x_{2},\ldots,x_{d})\coloneq\phi(x_{1})\times\phi(x_{2})\times\ldots\times\phi(x_{d})$ (8)
$\psi(x_{1},x_{2},\ldots,x_{d})\coloneq\psi(x_{1})\times\psi(x_{2})\times\ldots\times\psi(x_{d})$ (9)
Definition 4.
@@
Let $\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})=f(x^{\prime}_{1},x_{2},\ldots,x_{d})-f(x_{1},x_{2},\ldots,x_{d})$ such that $\delta_{x_{1}}$ is the difference operator in direction $x_{1}$ We say a function is mixed $\alpha$-Holder if the following condition is satisfied:
$\delta_{x_{d}}\delta_{x_{d-1}}\ldots\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})\leq C(|x^{\prime}_{1}-x_{1}|^{\alpha}|x^{\prime}_{2}-x_{2}|^{\alpha}\ldots|x^{\prime}_{d}-x_{d}|^{\alpha}),0<\alpha<\frac{1}{2}$ (10)
@@
We denote the space of functions satisfying this condition as $\Lambda^{\alpha}([0,1]^{d})$
@@
We now define the averaging operator acting on a function supported on a $d$-dimensional box. Let $I_{1}\times I_{2}\times\ldots\times I_{d}=[0,1]^{d}$. Along each interval we construct $j_{i}=0_{i},1_{i},2_{i},\ldots,N_{i}$ scales such that at scale $j$ one has $k_{i}=1_{i},2_{i},\ldots,2^{j_{i}}$ intervals. We denote the $k$th dyadic interval along scale $j$ in direction $i$ for the interval in the ith direction, $I_{i}$, as $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1))$ and $\chi^{j_{i}}_{k_{i}}(x_{i})$ as the characteristic function associated with interval $I^{j_{i}}_{k_{i}}$.
Definition 5.
We define the averaging operator acting on a function supported on a box as:
$P^{j_{1},j_{2},\ldots,j_{d}}(f)\coloneq\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}P^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f)$ (11)
$=\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}\frac{1}{|I^{j_{1}}_{k_{1}}|\times|I^{j_{2}}_{k_{2}}|\times\ldots\times|I^{j_{d}}_{k_{d}}|}(\int_{|I^{j_{1}}_{k_{1}}|}\int_{|I^{j_{2}}_{k_{2}}|}\ldots\int_{|I^{j_{d}}_{k_{d}}|}$ (12)
$\phi(y_{1},y_{2},\ldots,y_{d})f(y_{1},y_{2},\ldots,y_{d})dy_{1}dy_{2},\ldots dy_{d})\chi^{j_{1}}_{k_{1}}(x_{1})\times\chi^{j_{2}}_{k_{2}}(x_{2})\times\ldots\times\chi^{j_{d}}_{k_{d}}(x_{d})$ (13)
Definition 6.
Let the dyadic distance along direction $i$ be defined as:
$d_{i}(x^{\prime}_{i},x_{i})\coloneq\inf_{I_{i}}\{|I_{i}|:x^{\prime}_{i},x_{i}\in I_{i}\}$ (14)
2 d-tensor paraproducts
Theorem 2.
Suppose $A\in C^{d},f\in\Lambda^{\alpha}([0,1]^{d})$, then we obtain the following discrete $d$-tensor paraproduct decomposition for $A(f)$:
${A}(f)=\sum_{j_{1}=0_{1}}^{N_{1}}\sum_{j_{2}=0_{2}}^{N_{2}}\ldots\sum_{j_{d}=0_{d}}^{N_{d}}(\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}(f))+\Delta(A,f)$ (15)
$A(f)=\tilde{A}(f)+\Delta(A,f)$ (16)
where the residual $\Delta(A,f)$ is:
@@
$\Delta(A,f)\coloneq A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{d})$ (17)
$P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is the averaging operator acting on $f$ at scales $(j_{1},j_{2},\ldots,j_{d})$ and $\tilde{\mathbf{v}}^{\alpha}(f)$ is an operator acting on $f$ comprised of wavelets in $\alpha$ directions.
@@
Lemma 1.
Let $A\in C^{\infty}$ while $f\in\Lambda^{\alpha}([0,1]^{d})$. Define $Q^{j_{i}}$ as the finite difference operator between scales of the averaging operator:
@@
$=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$ (18)
@@
$=\sum_{k_{i}=0_{i}}^{2^{N_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$ (20)
@@
The fundamental theorem of calculus can be used to connect the identity established in lemma 3 to a quasilinearization of $A(f)$, but to do so we need to define a $d$-dimensional interpolation operator. This leads us to the next lemma.
Lemma 2.
@@
$\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{d}=0}^{t_{d}=1}\left.\right|_{t_{d-1}=0}^{t_{d-1}=1}\ldots\left.\right|_{t_{1}=0}^{t_{1}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$ (24)
@@
$\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1}=Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)$ (26)
@@
$\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{d}=0}^{t_{d}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$ (28)
@@
$\delta^{t}(f)\coloneq\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)$ (29)
such that lemma 2 holds
From the previous two lemmas and the fundamental theorem of calculus $A(f)$ has an identity which is a summation (with respect to the scaling parameters in all directions) over an integration of the power series comprised of the operators $(\mathbf{v}^{1}(f),\mathbf{v}^{2}(f),\ldots,\mathbf{v}^{d}(f))$ with respect to the interpolation parameters in all directions. First, we define the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ (setting each $\mathbf{v}(f)\coloneq\mathbf{v}^{m}$ to simplify notation) which come from differentiating $A(\delta^{t}(f))$. The elements in the sequence $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ are defined such that each power of $\mathbf{v}$ corresponds to the number of independent partial derivatives of $\delta^{t}(f)$ taken in each term of the series use to construct some $\mathbf{v}^{m}\in(\mathbf{v}^{1},\mathbf{v}^{2}\ldots,\mathbf{v}^{d})$:
$\mathbf{v}^{1}\coloneq\frac{\partial^{d}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (30)
$\mathbf{v}^{2}\coloneq\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\frac{\partial}{\partial t_{2}}(\delta^{t}(f))\frac{\partial}{\partial t_{1}}(\delta^{t}(f)))+$
$\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{d-1}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (31)
@@
$\mathbf{v}^{m}\coloneq\frac{\partial^{m}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
@@
$\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{m}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{d-m-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-m-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))\in[0,1]^{d}\to\mathbb{R}$ (32)

### v2

@@ Abstract.
We generalize the 22-tensor paraproduct decomposition result of [2] to $d$-tensors. In particular, we show that for $A\in C^{d}(\mathbb{R}),f\in\Lambda_{\alpha}([0,1]^{d})$, $A(f)$ can be approximated by $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=(\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f))$ with the residual $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=\tilde{A}_{(N_{i})_{i=1}^{d}}(f)-A(f)\in\Lambda_{2\alpha}([0,1]^{d})$. Our theoretical findings are supported by a computational example for d=3.
1. Introduction
We generalize the 2-tensor paraproduct decomposition developed in [2], which built on Bony’s seminal work, [1], to $d$-tensors for the situation $(f\in\Lambda_{\alpha}([0,1]^{d}),A\in C^{d}(\mathbb{R}))$; in particular, we prove the following theorem:
Theorem $1.$
Suppose $A\in C^{d}(\mathbb{R})$, $f\in\Lambda_{\alpha}([0,1]^{d}),0<\alpha<\frac{1}{2}$, then for the operator $T:f\to A(f)$ we can approximate $A(f)$ with
(1) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f)\in\Lambda_{\alpha}([0,1]^{d})$
such that the multiscale tensor paraproduct transforms $T:f\to A(f)$ to
(2) $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}:f\to\tilde{A}_{(N_{i})_{i=0}^{d}}(f)+\Delta_{(N_{i})_{i=0}^{d}}(A,f)$
@@ 1. Introduction
where $\Delta_{(N_{i})_{i=0}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{2})$ is the residual containing twice the regularity of $f$.
Furthermore, examinations of the organs of $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}$ leads to caricatures highlighted herein - namely, that (a) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$ is a taylor expansion of AA in terms of the derivatives of $A$ and scaling parameters of the wavelet operators and (b) the 3-tensor paraproduct decomposition.
2. Acknowledgments
The author wishes to thank Ronald R. Coifman for useful discussions.
3. Preliminaries
Included in this section are the relevant definitions needed to build $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}$:
Definition 3.1.
Dyadic distance along direction $i$:
(3) $d_{d}(x^{\prime}_{i},x_{i})=\inf_{j_{i},k_{i}}\begin{cases}|I^{j_{i}}_{k_{i}}|&x^{\prime}_{i},x_{i}\in I^{j_{i}}_{k_{i}}\\ 0,&x^{\prime}_{i}=x_{i}\end{cases}$
where $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1)]$ is the dyadic interval at scale $j_{i}$ at location $k_{i}$.
Definition 3.2.
The 1D scaling and wavelet functions have the usual definitions:
(4) $\phi(x)\coloneq\begin{cases}1&\text{if }x\in(0,1]\\ 0&\text{if }x\notin(0,1]\end{cases}\qquad\psi(x)\coloneq\begin{cases}1&\text{if }0<x\leq\frac{1}{2}\\ -1&\text{if }\frac{1}{2}<x\leq 1\end{cases}$
(5) $\phi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\phi(2^{-j}x-k),j,k\in\mathbb{N}\qquad\psi^{j}_{k}(x)\coloneq 2^{\frac{-j}{2}}\psi(2^{-j}x-k),j,k\in\mathbb{N}$
Definition $3.$
@@ 3. Preliminaries
(6) $\phi^{j}_{k}(x_{1},x_{2},\ldots,x_{d})\coloneq\phi^{j}_{k}(x_{1})\times\phi^{j}_{k}(x_{2})\times\ldots\times\phi^{j}_{k}(x_{d})$
(7) $\psi^{j}_{k}(x_{1},x_{2},\ldots,x_{d})\coloneq\psi^{j}_{k}(x_{1})\times\psi^{j}_{k}(x_{2})\times\ldots\times\psi^{j}_{k}(x_{d})$
Definition 3.4.
@@
Let $\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})=f(x^{\prime}_{1},x_{2},\ldots,x_{d})-f(x_{1},x_{2},\ldots,x_{d})$ such that $\delta_{x_{1}}$ is the difference operator in direction $x_{1}$ We say a function is mixed $\alpha$-Hölder if the following conditions are satisfied:
(8) $\delta_{x_{d}}\delta_{x_{d-1}}\ldots\delta_{x_{1}}f(x_{1},x_{2},\ldots,x_{d})\leq C(d_{d}(x^{\prime}_{1},x_{1})^{\alpha}d_{d}(x^{\prime}_{2},x_{2})^{\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{\alpha}),0<\alpha<\frac{1}{2}$
(9) $\delta_{x_{i}}f(x_{1},x_{2},\ldots,x_{d})\leq Cd_{d}(x^{\prime}_{i},x_{i})^{\alpha}\hskip 5.69046pt,\hskip 5.69046pti=1,\ldots,d$
@@
for $C,\alpha>0$. We denote the space of functions satisfying this condition as $\Lambda_{\alpha}([0,1]^{d})$
@@
We now define the averaging operator acting on a function supported on a $d$-dimensional box. Let $I_{1}\times I_{2}\times\ldots\times I_{d}=[0,1]^{d}$. Along each direction, $i$, we construct $j_{i}=0,1,2,\ldots,N$ scales such that at the $j_{i}$th scale one has $k_{i}=1,2,\ldots,2^{j}$ intervals. Let the dyadic interval along direction $i$ be denoted by: $I^{j_{i}}_{k_{i}}=(2^{j_{i}}k_{i},2^{j_{i}}(k_{i}+1)]$ and let $\chi^{j_{i}}_{k_{i}}(x_{i})$ as the characteristic function associated with interval $I^{j_{i}}_{k_{i}}$.
Definition 3.5.
Let the averaging operator, $P^{j_{1},j_{2},\ldots,j_{d}}$, be understood with the relation:
$P^{j_{1},j_{2},\ldots,j_{d}}(f)\coloneq\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}P^{j_{1},j_{2},\ldots,j_{d}}_{k_{1},k_{2},\ldots,k_{d}}(f)$
$=\sum_{k_{1}=1}^{2^{j_{1}}}\sum_{k_{2}=1}^{2^{j_{2}}}\ldots\sum_{k_{d}=1}^{2^{j_{d}}}\frac{1}{|I^{j_{1}}_{k_{1}}|\times|I^{j_{2}}_{k_{2}}|\times\ldots\times|I^{j_{d}}_{k_{d}}|}(\int_{|I^{j_{1}}_{k_{1}}|}\int_{|I^{j_{2}}_{k_{2}}|}\ldots\int_{|I^{j_{d}}_{k_{d}}|}$
(10) $\phi(y_{1},y_{2},\ldots,y_{d})f(y_{1},y_{2},\ldots,y_{d})dy_{1}dy_{2},\ldots dy_{d})\chi^{j_{1}}_{k_{1}}(x_{1})\times\chi^{j_{2}}_{k_{2}}(x_{2})\times\ldots\times\chi^{j_{d}}_{k_{d}}(x_{d})$
4. d-tensor paraproducts
Included in this section is a statement and proof of our main result:
Theorem 4.1.
Suppose $A\in C^{d}(\mathbb{R})$, $f\in\Lambda_{\alpha}([0,1]^{d}),0<\alpha<\frac{1}{2}$, then for the operator $T:f\to A(f)$ we can approximate $A(f)$ with
(11) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},\ldots,j_{d}=0}^{N_{1},\ldots,N_{d}}\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}(f)\in\Lambda_{\alpha}([0,1]^{d})$
such that the multiscale tensor paraproduct transforms $T:f\to A(f)$ to
(12) $\Pi^{(N_{i})_{i=1}^{d}}_{(A^{\beta})_{\beta=1}^{d}}:f\to\tilde{A}_{(N_{i})_{i=0}^{d}}(f)+\Delta_{(N_{i})_{i=0}^{d}}(A,f)$
@@
where $\Delta_{(N_{i})_{i=0}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{2})$ is the residual containing twice the regularity of $f$.
@@
Lemma 4.2.
Let $A\in C^{d}(\mathbb{R})$ and $f\in\Lambda_{\alpha}([0,1]^{d})$. Define $Q^{j_{i}}$ as the finite difference operator between scales of the averaging operator:
@@
(13) $=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$
@@
(15) $=\sum_{k_{i}=0_{i}}^{2^{N_{i}}}\frac{1}{|I^{{j_{i}}}_{k_{i}}|}\int_{|I^{{j_{i}}}_{k_{i}}|}\psi^{j_{i}}_{k_{i}}(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})f(x_{1},x_{2},\ldots,y_{i},\ldots,x_{d})dy_{i}\times\chi^{{j_{i}}}_{k_{i}}(x_{i})$
@@
The fundamental theorem of calculus can be used to connect the identity established in lemma 4.5 to a quasilinearization of $A(f)$, but to do so we need to define a $d$-dimensional interpolation operator. This leads us to the next lemma.
Lemma 4.3.
@@
(19) $\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{d}=0}^{t_{d}=1}\left.\right|_{t_{d-1}=0}^{t_{d-1}=1}\ldots\left.\right|_{t_{1}=0}^{t_{1}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
@@
(21) $\delta^{t_{k}}\delta^{t_{k-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{k}=0}^{t_{k}=1}=Q^{j_{k}}Q^{j_{k-1}}\ldots Q^{j_{1}}(f)$
@@
(23) $\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)\left.\right|_{t_{1}=0}^{t_{1}=1}\left.\right|_{t_{2}=0}^{t_{2}=1}\ldots\left.\right|_{t_{d}=0}^{t_{d}=1}=Q^{j_{d}}Q^{j_{d-1}}\ldots Q^{j_{1}}(f)$
@@
(24) $\delta^{t}(f)\coloneq\delta^{t_{d}}\delta^{t_{d-1}}\ldots\delta^{t_{1}}(f)$
such that lemma 4.3 holds
From the previous two lemmas and the fundamental theorem of calculus $A(f)$ has an identity which is a summation (with respect to the scaling parameters in all directions) over an integration of the power series comprised of the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d}$ with respect to the interpolation parameters in all directions. First, we define the operators $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ which come from differentiating $A(\delta^{t}(f))$. The elements in $(\mathbf{v}^{1},\mathbf{v}^{2},\ldots,\mathbf{v}^{d})$ are defined such that each power of $\mathbf{v}$ corresponds to the number of independent partial derivatives of $\delta^{t}(f)$ taken:
(25) $\mathbf{v}^{1}(f)\coloneq\frac{\partial^{d}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{1}}(\delta^{t}(f))$
$\mathbf{v}^{2}(f)\coloneq\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\frac{\partial}{\partial t_{2}}(\delta^{t}(f))\frac{\partial}{\partial t_{1}}(\delta^{t}(f)))+$
(26) $\frac{\partial^{d-2}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{3}}(\delta^{t}(f))\frac{\partial^{2}}{\partial t_{2}\partial t_{1}}(\delta^{t}(f))+\ldots+\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{d-1}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{1}}(\delta^{t}(f))$
@@
$\mathbf{v}^{m}(f)\coloneq\frac{\partial^{m}}{\partial t_{d}\partial t_{d-1}\ldots\partial t_{m}}(\delta^{t}(f))\frac{\partial}{\partial t_{m-1}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$
@@
(27) $\frac{\partial}{\partial t_{d}}(\delta^{t}(f))\frac{\partial^{m}}{\partial t_{d-1}\partial t_{d-2}\ldots\partial t_{d-m-1}}(\delta^{t}(f))\frac{\partial}{\partial t_{d-m-2}}(\delta^{t}(f))\ldots\frac{\partial}{\partial t_{1}}(\delta^{t}(f))$